文档下载 文档下载导航
设为首页 | 加入收藏
搜索 请输入内容:  
 导航当前位置: 文档下载 > 所有分类 > 高等教育 > 经济学 > 罗斯《公司理财》英文习题答案DOCchap012
侵权投诉

罗斯《公司理财》英文习题答案DOCchap012

b.

12.6 a.

b. c. d.

B-124

The mechanics of computing the means and the deviations were presented in an earlier chapter.

RT = 0.196 / 12 = 0.016333 RM = 0.236 / 12 = 0.019667 E(RT -RT) (RM -RM) = 0.038711 E(RM -RM)2 = 0.038588 = 0.038711 / 0.038588 = 1.0032

The beta of the average stock is 1. Mercantile’s beta is close to 1, indicating that its stock has average risk.

RM can have three values, 0.16, 0.18 or 0.20. The probability that RMtakes one of these values is the sum of the joint probabilities of the return pair that include the particular value of RM. For example, if RM is 0.16, RJ will be 0.16, 0.18 or 0.22. The probability that RM is 0.16 and RJ is 0.16 is 0.10. The probability that RM is 0.16 and RJ is 0.18 is 0.06. The probability that RM is 0.16 and RJ is 0.22 is 0.04. The probability that RM is 0.16 is, therefore, 0.10 + 0.06 + 0.04 = 0.20. The same procedure is used to calculate the probability that RM is 0.18 and the probability that RM is 0.20. Remember, the sum of the probability must be one.

RM

Probability 0.16 0.20 0.18 0.60 0.20 0.20 i.

R

M

= 0.16 (0.20) + 0.18 (0.60) + 0.20 (0.20) = 0.18

ii. 2M = (0.16 - 0.18) 2 (0.20) + (0.18 - 0.18) 2 (0.60) + (0.20 - 0.18) 2 (0.20)

= 0.00016

iii. M = 0.00016

1

2 = 0.01265

.18 .20 .20 .40 .22 .20 .24

.10

i. Ej = .16 (.10) + .18 (.20) + .20 (.40) + .22 (.20) + .24(.10) = .20 ii. j2 = (.16 - .20)2 (.10) + (.18 - .20)2 (.20) + (.20 - .20)2 (.40) + (.22 - .20)2 (.20) + (.24 - .20)2 (.10) = .00048

Answers to End-of-Chapter Problems

上一页第2页

你可能喜欢

  • 罗斯公司理财
  • 笔试题库
  • 10套计量经济学试卷(附答案)
  • 2010年考研政治真题及答案
  • 船舶操纵题库
  • 2010陕西行测真题及答案
  • 2010年6月四级真题答案

热门文档

相关文档

站点地图 | 文档上传 | 侵权投诉 | 手机版
本站所有资源均来自互联网,本站只负责收集和整理,均不承担任何法律责任,如有侵权等其它行为请联系我们.
文档下载 Copyright 2013 doc.guandang.net All Rights Reserved.  -  email